# Symbolic Higher-Order Analysis of Multivariate Time Series

### v1

@@ Acknowledgements.
A.C. and V.L. acknowledge support from the European Union - NextGenerationEU, GRINS project (grant E63-C22-0021-20006).
@@
We tested our method on artificial symbolic sequences containing known $2$- and $3$-motifs with tunable noise. To generate these sequences, we used an alphabet of $N$ symbols (integers from 00 to $N-1$).
@@
We end up with a random sentence composed of $r_{2}$ 2-words and $r_{3}$ 3-words, which represents the signal. The total length of the sentence is $L_{s}=r_{3}n3+r_{2}n_{2}$. We then add noise to the sentence, consisting of symbols randomly drawn from the alphabet (optionally including the special “space” character, associated with the number $-1$), which are inserted between adjacent words in the sentence. The level of noise is quantified by the noise-to-signal ratio $r_{ns}$, defined as the ratio between the total number of symbols in the noise, $L_{n}$, and the number of symbols in the signal, $L$:
@@
Figure S1: Distributions of the observed occurrences of unique tuples and motifs (i.e., significant tuples) in an artificial symbolic sequence created with our algorithm, for $2$-tuples (left) and $3$-tuples (right). These distributions refers to an artificial symbolic series created using an alphabet of $100$ symbols (plus the empty-space) and the following parameters: $n_{2}=478$, $n_{3}=499$ (starting from 500 randomly created 2-motifs and 500 3-motifs, with duplicates subsequently removed), $r_{2}=175$, $r_{3}=25$, $r_{ns}=10$, and $\gamma=-1$.
@@
Figure S2: Testing the method on an artificial sequence with noise-to-signal ratio $r_{ns}=10$ and $\gamma=-1$ (linear rank-frequency distribution). (a–b) Comparison between prior $\Pi(p)$ and posterior $P(p\mid\text{Data})$ distributions for a non-significant tuple and a significant one (i.e., a motif), respectively. (c–d) Histogram of the number of $2$- and $3$-tuples as a function of the Jensen-Shannon distance $d_{\text{JS}}$. A logarithmic scale is used for the $y$-axis in panel (d) to improve readability, given the larger number of possible $3$-tuples compared to motifs.
@@
Figure S3: Confusion matrices for the BJS-score in the classification of $2$-tuples (a) and $3$-tuples (b) in artificial data, using a significance threshold of $\text{BJS}^{\text{thr}}=0.6$. (c,d) Bar plots showing the corresponding performance metrics of the BJS-score compared to those of the z-score, computed on the same data with a threshold of $z^{\text{thr}}=3$. Our method, based on the BJS-score, consistently outperforms the z-score, particularly in the detection of higher-order motifs.
@@
From our analysis of ROC and PR curves on artificial sequences, we found that across all tested noise distributions (corresponding to several values of $\gamma$), the optimal $\text{BJS}^{\text{thr}}(l)$—that is, the value maximizing the F1 score—consistently falls between $0.5$ and $0.7$, for both $l=2$ and $l=3$. This provides a practical rule of thumb: this range is typically a good starting point for exploring threshold values and often yields good performance. This threshold also has a natural interpretation. The $\text{BJS}(\bm{s})$ measures the distance between prior and posterior distributions and takes values in the interval $[0,1]$, where 00 indicates identical distributions and $1$ indicates completely different ones. Hence, a threshold above the midpoint $0.5$ intuitively corresponds to distributions that are more different than similar.
@@
We tested this method on artificial sequences generated with different values of $\gamma$, and found that, although the threshold estimated using this procedure in general does not perfectly match the optimal $\text{BJS}^{\text{thr}}(l)$ (i.e., the one maximizing the F1 score), the typical range of both estimated and optimal thresholds is very similar—approximately between $0.5$ and $0.7$—thus confirming the reliability of this interval as a reasonable default choice.
@@
The z-score of a $l$-tuple $\bm{s}$ is defined as follow:
[…]
where $p_{\text{exp}}(\bm{s})$ is the expected probability for observing $\bm{s}$ taking into account the effect of lower-order correlations coming from tuples of length smaller than $l$ [43], defined in the main manuscript. For the ordered case we have:
$p_{\text{exp}}(\bm{s}=s_{1},s_{2})$ $=p_{\text{obs}}(s_{1})p_{\text{obs}}(s_{2})$ (S7)
@@
$p_{\text{exp}}(\bm{s}=s_{1},\dots,s_{l})$ $=\frac{p_{\text{obs}}(s_{1},\dots,s_{l-1})p_{\text{obs}}(s_{2},\dots,s_{l})}{p_{\text{obs}}(s_{2},\dots,s_{l-1})}$ (S8)
for $l>2$. We recall from the manuscript, that the observed probability of $\bm{s}$ is computed as:
$p_{\text{obs}}(\bm{s})=\frac{n_{\text{obs}}(\bm{s})}{S-l+1}$ (S9)
where $n_{\text{obs}}(\bm{s})$ represents the number of occurrences of $\bm{s}$ in the sequence of length $S$ and $S-l+1$ is the total number of $l$-tuples. For the application to the case of unordered tuples, all we have to do is to sum the results for the ordered case over $S_{p}(\bm{s})$, the set of unique permutations $\bm{\sigma}$ of the ordered tuple $\bm{s}=s_{1},\dots,s_{l}$, obtaining:
$p_{\text{exp}}^{\text{und}}(\bm{s})=\sum_{\bm{\sigma}\in S_{p}(\bm{s})}p_{\text{exp}}(\bm{\sigma})$ (S10)
@@
$n_{\text{obs}}^{\text{und}}(\bm{s})=\sum_{\bm{\sigma}\in S_{p}(\bm{s})}n_{\text{obs}}(\bm{\sigma})$ (S11)
In order to estimate the standard deviation $\text{std}_{\text{exp}}(\bm{s})$, we partition the original sequence in $m$ successive sub-sequences (namely, batches) of equal length. In each batch $i$ we compute $p_{\text{exp}}^{(i)}(\bm{s})$ and then we compute the standard deviation of $\bm{s}$ as:
$\text{std}_{\text{exp}}(\bm{s})\approx\sqrt{\frac{\langle p_{\text{exp}}^{(i)}(\bm{s})^{2}\rangle-\langle p_{\text{exp}}^{(i)}(\bm{s})\rangle^{2}}{m-1}}$ (S12)
[…]
@@
Data set $N$ $E_{2}$ $E_{3}$ $\text{BJS}^{\text{thr}}$
@@
Table S1: Basic properties of the unweighted hypergraphs obtained from multivariate time-series, respectively from single-cell neuronal activity [48], financial stocks [49, 50], and email exchanges [52]. Values of $\text{BJS}^{\text{thr}}=0.6,0.7$ have been adopted to threshold the weighted hypergraphs. A larger threshold means that fewer hyperedges on average satisfy the condition $\text{BJS}^{\text{thr}}\leq\text{BJS}(\bm{s})$, resulting in more sparse hypergraphs (i.e., with less hyperedges). For the stocks and the emails network we show the results ony for $\text{BJS}^{\text{thr}}=0.6$, as adopting a threshold of $0.7$ results in no higher-order hyperedges.
@@
We analyzed publicly available time series of neural activity with single-neuron resolution, provided by the Allen Institute. These datasets consist of multiple recording sessions performed on mice engaged in a visual task. For each mouse, approximately $500$ individual neurons were recorded during each session using Neuropixels technology. After filtering out neurons based on quality metrics—specifically retaining only those with an inter-spike interval (ISI) violation fraction $\text{isi}\leq 0.1$ and a signal-to-noise ratio $\text{snr}\geq 1.5$—we typically retained approximately 300300300300400400400400 neurons per session [48]. We applied our method to both the individual-neuron time series (and corresponding symbolic sequences), and an aggregated version in which neurons were grouped by their functional brain areas, as provided in the dataset metadata. Table S1 reports the number of nodes $N$, the number of $2$-motifs $E_{2}$, and the number of $3$-motifs $E_{3}$ in the unweighted hypergraphs obtained using significance thresholds $\text{BJS}^{\text{thr}}=0.6$ and $\text{BJS}^{\text{thr}}=0.7$. We present results for three different sessions of the “Brain Observatory” type [48], each corresponding to a different mouse. For each mouse, we analyzed all data available from the session, resulting in symbolic sequences of approximately $3\times 10^{7}$ symbols. In particular, “Brain 1” refers to session 773418906, “Brain 2” to session 791319847, and “Brain 3” to session 797828357 of the “Brain Observatory 1.1” dataset.
@@
We first converted each time series of closing prices into a binary time series as follows: if the change in closing price of a stock exceeded $2\%$ compared to the previous closing price, we assigned a value of $1$ in the corresponding binary time series (representing a spike); otherwise, we assigned a 00. We then applied the procedure described in the manuscript to map the binary multivariate time series of the $24$ stocks into a unique symbolic sequence of $25$ symbols (one for each stock, plus a special empty-space symbol). Given the daily timescale of this dataset, we inserted the empty symbol in the sequence whenever the time gap between two successive symbols exceeded $24$ hours. Finally, we extracted the $2$- and $3$-motifs from the symbolic sequence, as described in the manuscript. For this dataset, we repeated the entire procedure several times and observed minor fluctuations (on the order of ±$\pm 1$) in the number of motifs identified. This variability arises from the limited temporal resolution of the dataset, which causes many spikes (i.e., entries with value $1$ in the binary time series) to occur simultaneously. In such cases, to construct the symbolic time series, we must randomize the order of events that occur at the same time. This step introduces a degree of stochasticity in the mapping process, which in turn affects the final set of motifs. In all realizations of the procedure, we found that approximately $76\%$ of the pairwise motifs involved stocks belonging to the same category—an intuitively plausible result that supports the validity of our methodology. Regarding the $3$-motifs, our method identified as the most significant motifs the triplet (BAC, C, JPM), corresponding to the three banking stocks, and (COP, CVX, XOM), corresponding to the three energy stocks. Both motifs had a BJS-score of approximately $0.64\pm 0.02$.
We also considered the case of signed variations in stock prices by assigning each stock two distinct symbols: one representing a positive variation and the other a negative variation. Interestingly, in the case of signed changes, all the $2$-motifs—except one—involved stock price changes of the same sign ($36$ among symbols representing positive variations and $33$ among those representing negative variations). The only motif composed of symbols with opposite signs is (DOW$+$, DOW--), indicating that for the DOW stock, a variation in one direction is statistically associated with a subsequent correction in the opposite direction. In Table S1 we report the results relative to $\text{BJS}^{\text{thr}}=0.6$.

### v2

@@ Acknowledgements.
A.C. and V.L. acknowledge support from the European Union - NextGenerationEU, GRINS project (grant E63-C22-0021-20006). A.C and F.d.V.F. acknowledge support from the Agence Nationale de la Recherche under the France 2030 program (Grant ANR-23-IACL-0008).
@@
We tested our method on artificial symbolic sequences containing known $2$- and $3$-motifs with tunable noise. To generate these sequences, we used an alphabet of $N$ symbols (integers from 0 to N1N-1).
@@
We end up with a random sentence composed of $r_{2}$ 2-words and $r_{3}$ 3-words, which represents the signal. The total length of the sentence is $L_{s}=r_{3}n3+r_{2}n_{2}$. We then add noise to the sentence, consisting of symbols randomly drawn from the alphabet (optionally including the special “space” character, associated with the number 1-1), which are inserted between adjacent words in the sentence. The level of noise is quantified by the noise-to-signal ratio $r_{ns}$, defined as the ratio between the total number of symbols in the noise, $L_{n}$, and the number of symbols in the signal, $L$:
@@
Figure S1: Distributions of the observed occurrences of unique tuples and motifs (i.e., significant tuples) in an artificial symbolic sequence created with our algorithm, for 22-tuples (left) and 33-tuples (right). These distributions refers to an artificial symbolic series created using an alphabet of $100$ symbols (plus the empty-space) and the following parameters: $n_{2}=478$, $n_{3}=499$ (starting from 500 randomly created 2-motifs and 500 3-motifs, with duplicates subsequently removed), $r_{2}=175$, $r_{3}=25$, $r_{ns}=10$, and $\gamma=-1$.
@@
Figure S2: Testing the method on an artificial sequence with noise-to-signal ratio $r_{ns}=10$ and $\gamma=-1$ (linear rank-frequency distribution). (a–b) Comparison between prior $\Pi(p)$ and posterior $P(p\mid\text{Data})$ distributions for a non-significant tuple and a significant one (i.e., a motif), respectively. (c–d) Histogram of the number of $2$- and 33-tuples as a function of the Jensen-Shannon distance $d_{\text{JS}}$. A logarithmic scale is used for the $y$-axis in panel (d) to improve readability, given the larger number of possible 33-tuples compared to motifs.
@@
Figure S3: Confusion matrices for the BJS-score in the classification of $2$-tuples (a) and 33-tuples (b) in artificial data, using a significance threshold of $\text{BJS}^{\text{thr}}=0.6$. (c,d) Bar plots showing the corresponding performance metrics of the BJS-score compared to those of the z-score, computed on the same data with a threshold of $z^{\text{thr}}=3$. Our method, based on the BJS-score, consistently outperforms the z-score, particularly in the detection of higher-order motifs.
@@
From our analysis of ROC and PR curves on artificial sequences, we found that across all tested noise distributions (corresponding to several values of $\gamma$), the optimal $\text{BJS}^{\text{thr}}(l)$—that is, the value maximizing the F1 score—consistently falls between $0.5$ and $0.7$, for both $l=2$ and $l=3$. This provides a practical rule of thumb: this range is typically a good starting point for exploring threshold values and often yields good performance. This threshold also has a natural interpretation. The $\text{BJS}(\boldsymbol{s})$ measures the distance between prior and posterior distributions and takes values in the interval $[0,1]$, where 0 indicates identical distributions and $1$ indicates completely different ones. Hence, a threshold above the midpoint $0.5$ intuitively corresponds to distributions that are more different than similar.
@@
We tested this method on artificial sequences generated with different values of $\gamma$, and found that, although the threshold estimated using this procedure in general does not perfectly match the optimal $\text{BJS}^{\text{thr}}(l)$ (i.e., the one maximizing the F1 score), the typical range of both estimated and optimal thresholds is very similar—approximately between $0.5$ and 0.70.7—thus confirming the reliability of this interval as a reasonable default choice.
@@
The z-score of a $l$-tuple $\boldsymbol{s}$ is defined as follow:
$\text{z-score}(\boldsymbol{s})=\frac{p_{\text{obs}}(\boldsymbol{s})-p_{\text{exp}}(\boldsymbol{s})}{\text{std}_{\text{exp}}(\boldsymbol{s})}$ (S6)
where $p_{\text{exp}}(\boldsymbol{s})$ is the expected probability for observing $\boldsymbol{s}$ taking into account the effect of lower-order correlations coming from tuples of length smaller than $l$ [43], defined in the main manuscript. For the ordered case we have:
$p_{\text{exp}}(\boldsymbol{s}=s_{1},s_{2})$ $=p_{\text{obs}}(s_{1})p_{\text{obs}}(s_{2})$ (S7)
@@
$p_{\text{exp}}(\boldsymbol{s}=s_{1},\dots,s_{l})$ $=\frac{p_{\text{obs}}(s_{1},\dots,s_{l-1})p_{\text{obs}}(s_{2},\dots,s_{l})}{p_{\text{obs}}(s_{2},\dots,s_{l-1})}$ (S8)
for $l>2$. We recall from the manuscript, that the observed probability of $\boldsymbol{s}$ is computed as:
$p_{\text{obs}}(\boldsymbol{s})=\frac{n_{\text{obs}}(\boldsymbol{s})}{S-l+1}$ (S9)
where $n_{\text{obs}}(\boldsymbol{s})$ represents the number of occurrences of $\boldsymbol{s}$ in the sequence of length $S$ and Sl+1S-l+1 is the total number of $l$-tuples. For the application to the case of unordered tuples, all we have to do is to sum the results for the ordered case over $S_{p}(\boldsymbol{s})$, the set of unique permutations $\boldsymbol{\sigma}$ of the ordered tuple $\boldsymbol{s}=s_{1},\dots,s_{l}$, obtaining:
$p_{\text{exp}}^{\text{und}}(\boldsymbol{s})=\sum_{\boldsymbol{\sigma}\in S_{p}(\boldsymbol{s})}p_{\text{exp}}(\boldsymbol{\sigma})$ (S10)
@@
$n_{\text{obs}}^{\text{und}}(\boldsymbol{s})=\sum_{\boldsymbol{\sigma}\in S_{p}(\boldsymbol{s})}n_{\text{obs}}(\boldsymbol{\sigma})$ (S11)
In order to estimate the standard deviation $\text{std}_{\text{exp}}(\boldsymbol{s})$, we partition the original sequence in $m$ successive sub-sequences (namely, batches) of equal length. In each batch $i$ we compute $p_{\text{exp}}^{(i)}(\boldsymbol{s})$ and then we compute the standard deviation of $\boldsymbol{s}$ as:
$\text{std}_{\text{exp}}(\boldsymbol{s})\approx\sqrt{\frac{\langle p_{\text{exp}}^{(i)}(\boldsymbol{s})^{2}\rangle-\langle p_{\text{exp}}^{(i)}(\boldsymbol{s})\rangle^{2}}{m-1}}$ (S12)
where $\langle\cdot\rangle$ is the average over the $m$ batches. For our results, we partitioned the original sequence in $m=20$ batches.
@@
Data set NN $E_{2}$ $E_{3}$ $\text{BJS}^{\text{thr}}$
@@
Table S1: Basic properties of the unweighted hypergraphs obtained from multivariate time-series, respectively from single-cell neuronal activity [48], financial stocks [49, 50], and email exchanges [52]. Values of $\text{BJS}^{\text{thr}}=0.6,0.7$ have been adopted to threshold the weighted hypergraphs. A larger threshold means that fewer hyperedges on average satisfy the condition $\text{BJS}^{\text{thr}}\leq\text{BJS}(\boldsymbol{s})$, resulting in more sparse hypergraphs (i.e., with less hyperedges). For the stocks and the emails network we show the results ony for $\text{BJS}^{\text{thr}}=0.6$, as adopting a threshold of $0.7$ results in no higher-order hyperedges.
@@
We analyzed publicly available time series of neural activity with single-neuron resolution, provided by the Allen Institute. These datasets consist of multiple recording sessions performed on mice engaged in a visual task. For each mouse, approximately $500$ individual neurons were recorded during each session using Neuropixels technology. After filtering out neurons based on quality metrics—specifically retaining only those with an inter-spike interval (ISI) violation fraction $\text{isi}\leq 0.1$ and a signal-to-noise ratio $\text{snr}\geq 1.5$—we typically retained approximately 300300400400 neurons per session [48]. We applied our method to both the individual-neuron time series (and corresponding symbolic sequences), and an aggregated version in which neurons were grouped by their functional brain areas, as provided in the dataset metadata. Table S1 reports the number of nodes $N$, the number of $2$-motifs $E_{2}$, and the number of $3$-motifs $E_{3}$ in the unweighted hypergraphs obtained using significance thresholds $\text{BJS}^{\text{thr}}=0.6$ and $\text{BJS}^{\text{thr}}=0.7$. We present results for three different sessions of the “Brain Observatory” type [48], each corresponding to a different mouse. For each mouse, we analyzed all data available from the session, resulting in symbolic sequences of approximately $3\times 10^{7}$ symbols. In particular, “Brain 1” refers to session 773418906, “Brain 2” to session 791319847, and “Brain 3” to session 797828357 of the “Brain Observatory 1.1” dataset.
@@
We first converted each time series of closing prices into a binary time series as follows: if the change in closing price of a stock exceeded $2\%$ compared to the previous closing price, we assigned a value of $1$ in the corresponding binary time series (representing a spike); otherwise, we assigned a 0. We then applied the procedure described in the manuscript to map the binary multivariate time series of the $24$ stocks into a unique symbolic sequence of $25$ symbols (one for each stock, plus a special empty-space symbol). Given the daily timescale of this dataset, we inserted the empty symbol in the sequence whenever the time gap between two successive symbols exceeded $24$ hours. Finally, we extracted the $2$- and $3$-motifs from the symbolic sequence, as described in the manuscript. For this dataset, we repeated the entire procedure several times and observed minor fluctuations (on the order of $\pm 1$) in the number of motifs identified. This variability arises from the limited temporal resolution of the dataset, which causes many spikes (i.e., entries with value $1$ in the binary time series) to occur simultaneously. In such cases, to construct the symbolic time series, we must randomize the order of events that occur at the same time. This step introduces a degree of stochasticity in the mapping process, which in turn affects the final set of motifs. In all realizations of the procedure, we found that approximately $76\%$ of the pairwise motifs involved stocks belonging to the same category—an intuitively plausible result that supports the validity of our methodology. Regarding the $3$-motifs, our method identified as the most significant motifs the triplet (BAC, C, JPM), corresponding to the three banking stocks, and (COP, CVX, XOM), corresponding to the three energy stocks. Both motifs had a BJS-score of approximately $0.64\pm 0.02$.
We also considered the case of signed variations in stock prices by assigning each stock two distinct symbols: one representing a positive variation and the other a negative variation. Interestingly, in the case of signed changes, all the $2$-motifs—except one—involved stock price changes of the same sign ($36$ among symbols representing positive variations and $33$ among those representing negative variations). The only motif composed of symbols with opposite signs is (DOW++, DOW-), indicating that for the DOW stock, a variation in one direction is statistically associated with a subsequent correction in the opposite direction. In Table S1 we report the results relative to $\text{BJS}^{\text{thr}}=0.6$.